Assume that women's heights are normally distributed with a mean 64.5 in, and a standard deviation of 2.3 in.

a. if 1 woman is randomly selected, find the probability that her height is less than 65 in.

b. if 46 women are randomly selected, find the probability that they have a mean height less than 65 in.



1
Expert's answer
2021-03-16T08:17:09-0400

Solution:

Given, Population mean, "\mu=64.5\ in \quad"

And Population standard deviation, "\sigma=2.3\ in \quad"

(a) To find: "P(x<65\ in )=?"

The z -score is :

"\begin{array}{l}

z_{0} \quad=\frac{x_{0}-\mu}{\sigma} \\

\ \ \ \ \ \ \ \ \ =\frac{65-64.5}{2.3} =\frac{0.5}{2.3}
=0.22

\end{array}"

"\Rightarrow P(x<65\ in )=P(z<0.22)"

Finally, the required probability is

"\begin{array}{l}

P(x<65\ { in })=0.5871 =58.71 \%

\end{array}" [Using tables]

(b): Now, we must consider that we have a sample with size greater than one:

Given Population mean, "\mu=64.5\ in \quad"

And Population standard deviation, "\sigma=2.3\ in \quad"

Sample size, n=46

To find": P(\bar{x}<65\ in )=?"

The z -score is :

"\begin{aligned}

z_{0} &=\frac{\bar{x}_{2}-\mu}{\sigma / \sqrt{n}} \\

&=\frac{65-64.5}{2.3 / \sqrt{46}} \\

&=1.47

\end{aligned}"

"\Rightarrow P(\bar{x}<65\ in )=P(z<1.47)"

So, the required probability is 0.9292 [Using tables]